A ball of mass 240g is moving through the air at 20m/s with a gravitational potential energy of 70J. With what speed will the ball hit the ground?
1
Expert's answer
2018-05-06T08:07:08-0400
(mv^2)/2+E=(mV^2)/2
V=√(v^2+2E/m)=√(〖20〗^2+(2(70))/0.24)=31 m/s.